OpenStudy (hellokitty17):

Choose the correct simplification of the expression (c2)9. a.c18 b.c11 c.c512 d. c

pooja195 (pooja195):

\[\huge~(c^2)^9\]

pooja195 (pooja195):

Is that what you mean?

oregonduck (oregonduck):

similar example (c3)4=c3∗4 (c3)4=c12 note: this isn't the answer to your question as it's just a similar example

OpenStudy (hellokitty17):

I got b

oregonduck (oregonduck):

yup that is right

pooja195 (pooja195):

-.-

OpenStudy (mathstudent55):

When you raise a power to a power, you MULTIPLY powers.

OpenStudy (anonymous):

I mean I got B LOL sorry had trouble there

OpenStudy (e.mccormick):

OregonDuck, b is not right...

Nnesha (nnesha):

\[\huge\rm (x^m)^n = x^{m \times n}\]

OpenStudy (anonymous):

@e.mccormick i did the equation and i got stuck with A and B

pooja195 (pooja195):

What is \[\huge 9 \times 2=???\]

OpenStudy (mathstudent55):

@OregonDuck You gave a good example that was correct. \((c^3)^4 = c^{12}\) showing that the exponents are multiplied. Then the poster said the answer is C. c^11, which means the exponents were added, you said that was correct. Answer C. is incorrect bec the exponents need to be multiplied not added.

OpenStudy (e.mccormick):

\(\huge\rm x^mx^n = x^{m + n}\) is what some people seem to be thinking about, but this problem is not in that form.

OpenStudy (mathstudent55):

^above I meant answer B. c^11, not answer C.

OpenStudy (e.mccormick):

a) c18 b) c11 c) c512 d) c

oregonduck (oregonduck):

so a?

OpenStudy (mathstudent55):

Yes, @OregonDuck , A is correct.